OpenStudy (anonymous):

Multiply n(2n^2+n+1)

OpenStudy (anonymous):

This might help

OpenStudy (anonymous):

\[n\left(2n^2+n+1\right)=2n^3+n^2+n\]\[2n^2\cdot n+n\cdot n+1\cdot n=\ldots\,?\]